# VIPTR: A Vision Permutable Extractor for Fast and Efficient Scene Text Recognition

### v3

@@ Abstract
Scene Text Recognition (STR) is a challenging task that involves recognizing text within images of natural scenes. Although current state-of-the-art models for STR exhibit high performance, they typically suffer from low inference efficiency due to their reliance on hybrid architectures comprised of visual encoders and sequence decoders. In this work, we propose the VIsion Permutable extractor for fast and efficient scene Text Recognition (VIPTR), which achieves an impressive balance between high-performance and rapid inference speeds in the domain of STR. Specifically, VIPTR leverages a visual-semantic extractor with a pyramid structure, characterized by multiple self-attention layers, while eschewing the traditional sequence decoder. This design choice results in a lightweight and efficient model capable of handling inputs of varying sizes. Extensive experimental results on various standard datasets for both Chinese and English scene text recognition validate the superiority of VIPTR. Notably, the VIPTR-T (Tiny) variant delivers highly competitive accuracy on par with other lightweight models and achieves SOTA inference speeds. Meanwhile, the VIPTR-L (Large) variant attains greater recognition accuracy, while maintaining a low parameter count and favorable inference speed. Our proposed method provides a compelling solution for the STR challenge, which blends high accuracy with efficiency and greatly benefits real-world applications requiring fast and reliable text recognition. The code is publicly available at https://github.com/cxfyxl/VIPTR.
1 Introduction
@@
Text Recognition (TR) is crucial for multimodal document analysis, autonomous driving, office automation, and other applications Shi et al. (2016); Guo et al. (2022); Yang et al. (2020); Guo et al. (2023); Li et al. (2023a); Wu et al. (2023); Zhou et al. (2023). Recent deep learning advancements have spurred growth in Scene Text Recognition (STR), which aims to digitize text from images in natural settings for use in language processing tasks. Despite progress, STR still faces hurdles such as distorted text, varied fonts, shadows, and complex backgrounds. Efforts are ongoing to improve recognition precision, with an emphasis on creating models that are accurate, fast, and lightweight for versatile deployment.
Figure 1: Model Architecture Evolution for Scene Text Recognition.
@@
Scene text recognition initially relied on deep learning models like CRNNs Shi et al. (2016) that combined visual feature extraction with sequence processing, favored for their fast inference despite sensitivity to text disturbances. Subsequent models adopted encoder-decoder, autoregressive approaches, enhancing accuracy through contextual understanding but at the cost of slower word-by-word translation speeds. More recent advances involve visual semantic parsing frameworks that integrate semantic knowledge to improve recognition accuracy in specific contexts, which often require large models or intricate semantic systems, leading to efficiency compromise. Recent researchers 2021 (2020); Zhang et al. (2023) tried to develop simple and effective model structures by using the attention mechanism to aggregate the sequence features corresponding to the same character, in addition, align the CE loss and CTC loss to improve the convergence speed of the CTC loss. From the perspective of simplifying the model architecture, some models used only CNN Borisyuk et al. (2018) or ViT Atienza (2021) for identification, which is efficient in inference but has a large gap in accuracy compared to SOTA methods. Inspired by the pyramid structure of Swin Transformer Liu et al. (2021), the SVTR model Du et al. (2022), designed for single visual modality input, is proposed to simultaneously process the implicit local visual information in text images and unsupervised mining of global contextual semantic information. SVTR has four structural variants with different capacities.
@@
The STR models based on vision Transformer set a typical example for fast and effective STR schemes for a single visual modality. On one hand, it practices the effectiveness of replacing convolution with a self-attention mechanism to realize a new visual feature extractor. On the other hand, it reveals the importance of modeling global dependencies of lateral symbol sequences in text image recognition tasks, which even replace bidirectional LSTMs. Despite the outstanding performance of current ViT-based STR, it is clear that it still has several shortcomings: (1) The computation of self-attention relies on absolute position encoding and fixed-size mask operation, which limits the length of the input image; (2) In each stage, only one type of global and local relations is mined separately, and the distance between the steps of global dependencies is limited to save reasoning time, so the recognition effect of long text images is not good; (3) The actual inference time consumption of the model in general scenarios is still larger than that of CRNN with the same parameter magnitude.
@@
To further enhance the vision model and meticulously control computational complexity to achieve more efficient STR while maintaining high accuracy, we propose VIPTR—a vision-based approach for a faster, more accurate, and cross-language universal STR model. Inspired by Cross-Shaped Windows Vision Transformer (CSwin) Dong et al. (2022) and other visual models based on enhanced self-attention mechanisms, VIPTR first divides the text image evenly into multiple 2D patches, each patch contains part of the characters or background. This way of dividing patches facilitates the simultaneous extraction of local features of characters and the capture of potential contextual semantic dependencies between different characters. In each token mixing attention module, we try to use conditional positional embedding (CPE) Chu et al. (2021) combined with Manhattan self-attention (MaSA) Fan et al. (2023) and CSWin Attention to explore the local information of characters, and use Overlapping Spatial Reduction Attention (OSRA) Lou et al. (2023) module to model global dependencies. Specifically, along the horizontal direction Model the global dependency between characters in the axial direction, model the global relationship from top to bottom of the same character along the vertical axis, and align to achieve weight sharing. Compared with SWin, this method significantly reduces the time complexity and does not require the use of a fixed-size mask group to limit the length of the input image. At the same time, we also use a pyramid structure for multi-stage local and global information hybrid modeling, which greatly improves the feature extraction capabilities of text images. To sum up, VIPTR can surpass the recognition effect and inference efficiency of SVTR at a smaller model size.
The main contributions of this work can be summarized as.
@@
  • We verified that the single-vision model based on the self-attention mechanism can still achieve comparable accuracy to the high-level vision-language model in the scene text recognition task. At the same time, the use of sparse operators and different attention combinations can accelerate the calculation of the self-attention mechanism, and achieve the balance between performance and speed advantages that STR has been pursuing.
@@
  • We proposed VIPTR, a feature extraction module tailored for parsing text. The module can adapt to and accurately recognize the text image input of any length and has the versatility of cross-language recognition, which has a good application prospect.
@@
  • We verified the superiority of VIPTR on cross-language benchmark datasets and manually annotated industrial application datasets. Among them, VIPTRv1-L achieved SOTA in recognizing Chinese and English scene text. VIPTRv2-T achieves the most efficient inference while ensuring accuracy. It is inferred on an NVIDIA V100 GPU with a parameter of 5.1M and consumes an average of 3.3ms per text image.
2 Related Work
@@
Figure 2: Overall architecture of the proposed VIPTR. It is a four-stage network with three-stage height progressively decreasing. In each stage, a series of attention-mixing blocks are carried out and followed by a subsampling or combining operation. In particular, VIPTR designs two visual feature fusion modes in vision permutation blocks: series and parallel. At last, the recognition is conducted by the CTC decoder.
CNN-Based Vision Modules
@@
Convolutional neural networks (CNN) serve as pivotal tools for extracting visual information in previous tasks. VGG Simonyan and Zisserman (2014) demonstrated that a series of 3x3 convolutions is sufficient to train a high-performance model. Later, ResNet He et al. (2016) introduced residual connections along with the batch normalization (BN) layer, which enabled the training of exceedingly deep neural networks and further improved performance. In the realm of STR, the CNN-based vision model transforms a pyramid structure capable of processing local features across multiple stages. This transformation effectively compresses the two-dimensional image input into a one-dimensional feature sequence, allowing for decoding the sequence without reliance on recurrent neural networks (RNN). High recognition accuracy can also be obtained, which proves the effectiveness of locality bias extracted by CNN. We believe that the experience gained in CNN-based vision models, including the pyramid structure for extracting locality bias through small kernel convolutions, residual connections, and step-by-step downsampling, remains valid in current research. Therefore, these techniques continue to be utilized in our proposed STR architecture.
Transformer-Based Vision Modules
@@
ViT Dosovitskiy et al. (2020) is the first work to introduce Transformer Vaswani et al. (2017); Yang et al. (2022); Touvron et al. (2021) into the construction backbone of vision tasks. Experiments demonstrate that, with sufficient training data, ViT outperforms traditional CNNs in vision tasks. DeiT Touvron et al. (2021) introduces a new training strategy, enabling ViT to achieve superior performance on small-scale training sets such as ImageNet-1K. These methods demonstrate the efficacy of global dependency modeling through self-attention mechanisms. Later, the pyramid structure also enhances the performance of Transformer-based models across various vision tasks. Based on the pyramid structure, Swin Transformer proposes to use self-attention within each local window. This can be regarded as a use of locality bias. Vision Transformers are integrated into STR tasks as visual feature encoders, leading to a significant improvement in accuracy. ViTSTR Atienza (2021), visionLAN Wang et al. (2021b), and MaskOCR Lyu et al. (2022) have all come close to the SOTA effect. SVTR-T both ensures high performance and creates the fastest recognition efficiency. Although the self-attention mechanism has greatly improved the effect of text recognition, its computational complexity is quadratic compared to the number of input tokens. As the resolution of the input image increases, the inference time consumption also increases quadratically.
Efficient Self-Attention Based Vision Modules
@@
A challenging problem in the design of Transformer is that the computational cost of global self-attention is very expensive Vaswani et al. (2017); Yang et al. (2021, 2023), while local self-attention usually limits the interaction field of each token, resulting in insufficient information mining. The CSWin Self-Attention mechanism uses parallel computing to form the self-attention of the horizontal and vertical stripes of the cross-shaped window, achieving powerful feature modeling capabilities while limiting computing costs. Decomposed MaSA achieves efficient and significant feature representation by defining a spatial attention attenuation matrix and its decomposition paradigm based on two-dimensional Manhattan distance, effectively layering local information on a global scale. Spatial reduction attention (SRA) Wang et al. (2021a) effectively extracts global information by utilizing sparse token-region dependencies, and OSRA introduces overlapping spatial reduction (OSR) Lou et al. (2023) into SRA to achieve better performance by using larger and overlapping patches.
3 Method
3.1 Design Guidelines
@@
    Explicitly inject character features into the front of the network in the form of locality bias.
  2. 2.
    Design a new permutation method to effectively fuse the character component features of the Local and Global Attention modules in the middle layer.
  3. 3.
    Implicitly handle image aspect and width resolution in the network initialization structure, canceling or replacing all modules that depend on the input size.
3.2 Overall Architecture
@@
The schematic representation of the proposed VIPTR is depicted in Figure 2. It is a four-stage network with three height progressively decreased layers. Given an image text of size $H\times W\times 3$, it undergoes an initial transformation into $\frac{H}{4}\times\frac{W}{4}\times C_{0}$ patches through a two-layer CNN-based patch embedding. Since each patch is associated with a text character or part of the background in the image, these patches are called character components. Subsequently, four stages of feature extraction are performed at different scales, each stage consists of a series of self-attention-based hybrid blocks, followed by downsampling or global average pooling operations along the height dimension. Both Local and Global Attention blocks are employed to concentrate on character feature extraction and capture dependencies between character components. The extractor serves to characterize character component features at different distances and multi-scales and the contextual dependencies between character components, and a representation called $B$ with a size of $1\times\frac{W}{4}\times C_{3}$ is generated, which captures multi-granularity character features. Finally, the CTC decoder is used to perform linear prediction and deduplication on the character sequence to obtain the recognition result.
3.3 Patch Embedding
@@
The first step of the vision Transformer is to cut the image into several equal-sized patches, and then hand them over to the subsequent self-attention module for processing in the form of several sets of sequences. In this article, for an image text, first convert it from $X\in R^{H\times W\times 3}$ into several character components $B_{0}B_{1}\in R^{\frac{H}{4}\times\frac{W}{4}\times C_{0}}$. Alternatively, we implement the patch embedding by using two consecutive $3\times 3$ convolutions with stride $2$ and BNs. This scheme Du et al. (2022) gradually increases feature dimensions, which is beneficial to feature fusion.
Figure 3: Illustration of (a) Cross-Shaped Windows Self-Attention (CSWin), (b) Decomposed Manhattan Self-Attention (D-MaSA), (c) Multi-Head Self-Attention (MHSA), and (d) Overlapping Spatial Reduction Attention (OSRA).
3.4 Character and String Modeling

### v5

# SVIPTR: Fast and Efficient Scene Text Recognition with Vision Permutable Extractor

@@ Abstract.
Scene Text Recognition (STR) is an important and challenging upstream task for building structured information databases, which involves recognizing text within images of natural scenes. Although current state-of-the-art (SOTA) models for STR exhibit high performance, they typically suffer from low inference efficiency due to their reliance on hybrid architectures comprised of visual encoders and sequence decoders. In this work, we propose a VIsion Permutable extractor for fast and efficient Scene Text Recognition (SVIPTR), which achieves an impressive balance between high performance and rapid inference speeds in the domain of STR. Specifically, SVIPTR leverages a visual-semantic extractor with a pyramid structure, characterized by the Permutation and combination of local and global self-attention layers. This design results in a lightweight and efficient model and its inference is insensitive to input length. Extensive experimental results on various standard datasets for both Chinese and English scene text recognition validate the superiority of SVIPTR. Notably, the SVIPTR-T (Tiny) variant delivers highly competitive accuracy on par with other lightweight models and achieves SOTA inference speeds. Meanwhile, the SVIPTR-L (Large) attains SOTA accuracy in single-encoder-type models, while maintaining a low parameter count and favorable inference speed. Our proposed method provides a compelling solution for the STR challenge, which greatly benefits real-world applications requiring fast and efficient STR. The code is publicly available at https://github.com/cxfyxl/VIPTR.
Scene Text Recognition, Attention mechanism, Vision Transformer, Visual-semantic analysis, Length-Insensitive
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679618isbn: 979-8-4007-0436-9/24/10ccs: Computing methodologies Object recognitionccs: Computing methodologies Neural networksccs: Computing methodologies Scene understandingccs: Computing methodologies Image representations
1. Introduction
@@
Text Recognition (TR) is crucial for multimodal document analysis and building structured databases, autonomous driving, office automation, and other applications (Shi et al., 2016; Guo et al., 2022; Yang et al., 2020; Guo et al., 2023; Li et al., 2023a; Wu et al., 2023; Zhou et al., 2023). Recent deep learning advancements have spurred growth in Scene Text Recognition (STR), which aims to digitize text from images in natural settings for use in language processing tasks. Despite progress, STR still faces hurdles such as distorted text, varied fonts, shadows, and complex backgrounds. Efforts are ongoing to improve recognition precision, with an emphasis on creating models that are accurate, fast, and lightweight for versatile deployment.
@@
STR initially relied on deep learning models like CRNNs (Shi et al., 2016) that combined visual feature extraction with sequence processing (as shown in Figure 1(a)), favored for their fast inference despite sensitivity to text disturbances. Subsequent models adopted encoder-decoder, autoregressive approaches (as shown in Figure 1(b)), enhancing accuracy through contextual understanding but at the cost of slower word-by-word translation speeds. More recent advances involve visual semantic parsing frameworks (as shown in Figure 1(c)) that integrate semantic knowledge to improve recognition accuracy in specific contexts, which often require large models or intricate semantic systems, leading to efficiency compromise.
Recent researchers (Hu et al., 2020; Zhang et al., 2023) tried to develop simple and effective model structures by using the attention mechanism to aggregate the sequence features corresponding to the same character, or regularising the extracted sequence frame by frame for supervised learning of individual characters. These strategies are used to align the Cross-Entropy (CE) loss and Connectionist Temporal Classification (CTC) loss, on the one hand, the convergence speed of CTC loss is improved without affecting inference speed, and on the other hand, the character and sequence features are strengthened. From the perspective of simplifying the model architecture, some models used only CNN (Borisyuk et al., 2018) or ViT (Atienza, 2021) for identification, which is efficient in inference but has a large gap in accuracy compared to SOTA methods. Recent researchers (Wang et al., 2022; Cheng et al., 2023; Fujitake, 2024) have proposed a series of innovative optimization methods for feature sequence decoders, simplifying the encoder-decoder architecture and effectively promoting a balance between high performance and inference efficiency. In addition, inspired by the pyramid structure of Swin Transformer (Liu et al., 2021), the SVTR model (Du et al., 2022), designed for single visual modality input, is proposed to simultaneously process the implicit local visual information in text images and unsupervised mining of global contextual semantic information.
@@
The STR models based on vision Transformer set a typical example for fast and effective STR schemes for a single visual modality. On one hand, it practices the effectiveness of replacing convolution with a self-attention mechanism to realize a new visual feature extractor. On the other hand, it reveals the importance of modeling global dependencies of lateral symbol sequences in text image recognition tasks, which even replace bidirectional LSTMs (Sun et al., 2022) and sequence decoders. Despite the outstanding performance of current ViT-based STR, it is clear that it still has several shortcomings: (1) The computation of self-attention relies on absolute position encoding and fixed-size mask operation, which limits the length of the input image; (2) In each stage, only one type of global and local relations is mined separately, and the distance between the steps of global dependencies is limited to save reasoning time, so the recognition effect of long text images is not good; (3) The Resource occupancy during training and CPU inference time consumption of models in general scenarios are still larger than that of CRNN with the same parameter magnitude.
Figure 1. Model Architecture Evolution for Scene Text Recognition.
[…]
Model Architecture Evolution for Scene Text Recognition.
@@
To further enhance the vision model and meticulously control computational complexity to achieve more efficient STR while maintaining high accuracy, we propose SVIPTR—a visual-semantic-based approach for a faster, more accurate, and cross-language universal STR model. Inspired by Cross-Shaped Windows Vision Transformer (CSwin) (Dong et al., 2022) and other visual models based on enhanced self-attention mechanisms, SVIPTR first divides the text image evenly into multiple 2D patches, each patch contains part of the characters or background. This way of dividing patches facilitates the simultaneous extraction of local features of characters and the capture of potential contextual semantic dependencies between different characters. In each token mixing attention module, we try to use conditional positional embedding (CPE) (Chu et al., 2021) combined with Manhattan self-attention (MaSA) (Fan et al., 2023) and CSWin Attention to explore the local information of characters, and use Overlapping Spatial Reduction Attention (OSRA) (Lou et al., 2023) module to model global dependencies. Specifically, along the horizontal direction Model the global dependency between characters in the axial direction, model the global relationship from top to bottom of the same character along the vertical axis, and align to achieve weight sharing. Compared with SWin, this method significantly reduces the time complexity and does not require the use of a fixed-size mask group to limit the length of the input image. At the same time, we also use a pyramid structure for multi-stage local and global information hybrid modeling, which greatly improves the feature extraction capabilities of text images. To sum up, SVIPTR can ensure that the recognition effect and inference efficiency exceed the advanced vision Transformer-based recognition model at a smaller model scale.
The main contributions of this work can be summarized as:
@@
  • We verify that the use of sparse operators and different Permutations and combinations of self-attention mechanisms can accelerate the calculation of Vision Transformer, and the model built based on these mechanisms can still achieve comparable accuracy to high-level vision-language models in STR tasks. Integrated, achieving the balance between performance and speed advantages that STR has been pursuing.
@@
  • We proposed SVIPTR, a visual-semantic feature extraction model tailored for parsing image text. The model can accurately recognize cross-language image text input and is insensitive to the length of image input, which has a good application prospect.
@@
  • We verified the superiority of SVIPTR on cross-language benchmark datasets and manually annotated industrial application datasets. Among them, SVIPTRv1-L achieved higher accuracy than other ViT-based encoder models on both Chinese and English STR. SVIPTRv1-L combined with BiLSTM decoder achieves the SOTA level on STR benchmarks. SVIPTRv2-T achieves the most efficient inference while ensuring accuracy. It is inferred on an NVIDIA V100 GPU with a parameter of 5.1M and consumes an average of 3.3ms per text image.
2. Related Work
@@
Figure 2. Overall architecture of the proposed SVIPTR. It is a four-stage network with three-stage height progressively decreasing. In each stage, a series of attention-mixing blocks are carried out and followed by a subsampling or combining operation. In particular, SVIPTR designs two visual feature fusion modes in vision permutation blocks: series and parallel. At last, the recognition is conducted by the CTC decoder.
[…]
Overall architecture of the proposed SVIPTR. It is a four-stage network with three-stage height progressively decreasing. In each stage, a series of attention-mixing blocks are carried out and followed by a subsampling or combining operation. In particular, SVIPTR designs two visual feature fusion modes in vision permutation blocks: series and parallel. At last, the recognition is conducted by the CTC decoder.
2.1. CNN-Based Vision Modules.
@@
Convolutional neural networks (CNN) serve as pivotal tools for extracting visual information in previous tasks. VGG (Simonyan and Zisserman, 2014) demonstrated that a series of 3x3 convolutions is sufficient to train a high-performance model. Later, ResNet (He et al., 2016) introduced residual connections along with the batch normalization (BN) layer, which enabled the training of exceedingly deep neural networks and further improved performance. In the realm of STR, the CNN-based vision model transforms a pyramid structure capable of processing local features across multiple stages. This transformation effectively compresses the two-dimensional image input into a one-dimensional feature sequence, allowing for decoding the sequence without reliance on recurrent neural networks (RNN). High recognition accuracy can also be obtained, which proves the effectiveness of locality bias extracted by CNN. We believe that the experience gained in CNN-based vision models, including the pyramid structure for extracting locality bias through small kernel convolutions, residual connections, and step-by-step downsampling, remains valid in current research. Therefore, these techniques continue to be utilized in our proposed STR architecture.
2.2. Transformer-Based Vision Modules.
@@
ViT (Dosovitskiy et al., 2020) is the first work to introduce Transformer (Vaswani et al., 2017; Yang et al., 2022; Touvron et al., 2021) into the construction backbone of vision tasks. Experiments demonstrate that, with sufficient training data, ViT outperforms traditional CNNs in vision tasks. DeiT (Touvron et al., 2021) introduces a new training strategy, enabling ViT to achieve superior performance compared to CNNs on small-scale training sets such as ImageNet-1K. These methods demonstrate the efficacy of global dependency modeling through self-attention mechanisms. Later, the pyramid structure also enhances the performance of Transformer-based models across various vision tasks. Based on the pyramid structure, Swin Transformer proposes to use self-attention within each local window. This can be regarded as a use of locality bias. Vision Transformers are integrated into STR tasks as visual feature encoders, leading to a significant improvement in accuracy. VisionLAN (Wang et al., 2021a), SVTR, MGP-STR (Wang et al., 2022) and MaskOCR (Lyu et al., 2022) have all come close to the SOTA effect.
2.3. Sparse Self-attention and Spatial Modeling.
@@
A challenging problem in the design of Transformer is that the computational cost of global self-attention is very expensive (Vaswani et al., 2017; Yang et al., 2021, 2023), while local self-attention usually limits the interaction field of each token, resulting in insufficient information mining. The CSWin Self-Attention mechanism uses parallel computing to form the self-attention of the horizontal and vertical stripes of the cross-shaped window, achieving powerful feature modeling capabilities while limiting computing costs. Decomposed MaSA achieves efficient and significant feature representation by defining a spatial attention attenuation matrix and its decomposition paradigm based on two-dimensional Manhattan distance, effectively layering local information on a global scale. Spatial reduction attention (SRA) (Wang et al., 2021b) effectively extracts global information by utilizing sparse token-region dependencies, and OSRA introduces overlapping spatial reduction (OSR) (Lou et al., 2023) into SRA to achieve better performance by using larger and overlapping patches.
3. Method
3.1. Design Guidelines
@@
    Explicitly vectorizing the character features in the form of segmented chunks and injecting them into the front of the network.
  2. (2)
    Design permutation and combination modes of self-attention mechanism to effectively fuse visual and semantic information of characters embeddings in the middle layer.
  3. (3)
    Implicitly handling the length information of input images during network initialization and designing shape-insensitive self-attention mechanisms to replace all modules that depend on the input size.
3.2. Overall Architecture
@@
The schematic representation of the proposed SVIPTR is depicted in Figure 2. It is a four-stage network with three height progressively decreased layers. Given an image text of size $H\times W\times 3$, it undergoes an initial transformation into $\frac{H}{4}\times\frac{W}{4}\times C_{0}$ patches through a two-layer CNN-based patch embedding. Since each patch is associated with a text character or part of the background in the image, these patches are called character components. Subsequently, four stages of feature extraction are performed at different scales, each stage consists of a series of self-attention-based hybrid blocks, followed by dimension reduction in half using convolutional or global average pooling operations along the height dimension. Finally, the extractor generated a representation called $B$ with a size of $1\times\frac{W}{4}\times C_{3}$, which captures multi-granularity characters’ sequence features. At the same time, the CTC decoder is used to linearly predict $B$ to obtain the recognition result.
3.3. Patch Embedding
@@
The first step of the vision Transformer is to cut the image into several equal-sized patches, and then hand them over to the subsequent self-attention module for processing in the form of several sets of sequences. In this article, for an image text, first convert it from $X\in R^{H\times W\times 3}$ into several character components $B_{0}B_{1}\in R^{\frac{H}{4}\times\frac{W}{4}\times C_{0}}$. Alternatively, we implement the patch embedding by using two consecutive $3\times 3$ convolutions with stride $2$ and BNs. This scheme (Du et al., 2022) gradually increases feature dimensions, which is beneficial to feature fusion.
3.4. Character and String Modeling